OpenStudy (anonymous):

Can someone help me? Replace (...) by x> or equal or x< or equal so that the inequality will be true for any x: a) x^2-16x+64...0 b) 64+16x+x^2...0 c) -x^2-4x-4...0 d) -x^2+18x-81

OpenStudy (anonymous):

I would like to help but I am confused by what you mean in this problem

OpenStudy (anonymous):

x^2-16x+64x>=0 ??? x^2-16x+64x<=0 ??? These are your choices?

OpenStudy (anonymous):

yea

OpenStudy (kira_yamato):

x^2 - 16x + 64 = (x - 8)^2 ≥ 0 64 + 16x + x^2 = (x - 8)^2 ≥ 0 - x^2 - 4x - 4 = - (x^2 + 4x + 4) = - (x + 2)^2 ≤ 0 - x^2 + 18x - 81 = - (x^2 - 18x + 81) = - (x-9)^2 ≤ 0